Evaluate each function rule for the given value.
a.) y=.5^x for x=3
b.) f(x)=4.3^x for x=-2

Answer by ikleyn(52754)

.

(a)  0.5^3 = 0.5*0.5*0.5 = 0.125.



(b)  4.3^(-2) = use your calculator.